# A Fast Parallel Approach for Neighborhood-based Link Prediction by Disregarding Large Hubs

### v3

@@ 1. Introduction
Most real world networks are incomplete (Kim and Leskovec, 2011; Wang et al., 2014). These networks lie somewhere in the range of a deterministic and a purely random structure, and are thus partially predictable (Lü et al., 2015). Link prediction is the problem of identifying potentially missing/undiscovered connections in such networks (Marchette and Priebe, 2008; Kim and Leskovec, 2011), or even forecasting future connections (Bringmann et al., 2010; Juszczyszyn et al., 2011), by examining the current network structure. This is useful in various applications, such as recommending items for online purchase (Akcora et al., 2011), helping people to find potential collaborators (Mori et al., 2012; Tang et al., 2012), assessing trustworthiness of individuals (Alnumay et al., 2019), uncovering criminal activities and individuals (Berlusconi et al., 2016; Lim et al., 2019), and predicting new protein-protein interactions or generating hypotheses (Cannistraci et al., 2013; Nasiri et al., 2021).
@@
Similarity measures are frequently employed to predict the likelihood of missing or future links between unconnected nodes in a network (Wang et al., 2014; Arrar et al., 2023). The principle is straightforward: higher similarity indicates a greater likelihood of connection (Wang et al., 2014). The choice of metric depends on the network’s characteristics, with no single metric dominating across different datasets (Arrar et al., 2023; Zhou, 2021). Local / neighborhood-based similarity metrics such as Common Neighbors, Jaccard Coefficient, Sørensen Index, Salton Cosine similarity, Hub Promoted, Hub Depressed, Leicht-Holme-Nerman, Adamic-Adar, and Resource Allocation, which are based of are based on neighborhood information within a path distance of two, remain popular (Arrar et al., 2023; Wang et al., 2014). This is due to their simplicity, interpretability (Pai et al., 2019; Barbieri et al., 2014), computational efficiency (Garcia Gasulla and Cortés García, 2014), and the ability to capture underlying structural patterns.
@@
However, many studies (Gatadi and Swarupa Rani, 2023; Saifi et al., 2023; Benhidour et al., 2022; Mumin et al., 2022; Rafiee et al., 2020; Guo et al., 2019; Yang et al., 2015b; Papadimitriou et al., 2012; Wang et al., 2019) and network analysis software (Staudt et al., 2016; Csardi et al., 2006) use a baseline approach for link prediction, computing unnecessary similarity scores for all non-connected node pairs. Further, early studies often evaluate a limited number of algorithms on small networks — this can result in misleading conclusions (Zhou, 2021; Zhou et al., 2021). Further, much existing research does not address link prediction for large networks with close to a billion edges (Muscoloni et al., 2022; Mumin et al., 2022; Nasiri et al., 2021; Xian et al., 2021; Ghasemian et al., 2020; Mara et al., 2020; Wang et al., 2019; Xu et al., 2019; Mohan et al., 2017; Cui et al., 2016; Garcia Gasulla and Cortés García, 2014; Papadimitriou et al., 2012). As the collection of data, represented as graphs, reach unprecedented levels, it becomes necessary to design efficient parallel algorithms for link prediction on such graphs. While link prediction algorithms are often pleasingly parallel, most studies do not address the design of suitable data structures for efficient computation of scores.
@@
Further, the link prediction problem faces significant imbalance, with the number of known present links often several order of magnitude less than known absent links. This imbalance hinders the effectiveness of many link prediction methods, particularly on large networks (Wang et al., 2014; Garcia Gasulla and Cortés García, 2014). Thus, heuristics are needed to minimize the computation needed, without sacrificing on quality. Quality assessment measures for link prediction include precision, recall, F1 score, accuracy, and Area Under the Receiver Curve (AUC). While AUC is commonly used (Arrar et al., 2023), it may provide misleading results (Yang et al., 2015a; Lichtnwalter and Chawla, 2012), leading to our focus on F1 score.
@@ 2. Related work
While effective in capturing nonlinear relationships for accurate predictions in complex networks, embedding-based methods face challenges with high-centrality nodes. Matrix factorization methods often struggle with complex network structures, require substantial computational resources and pose overfitting risks (Arrar et al., 2023; Martínez et al., 2016). Machine learning based approaches have their own set of drawbacks. These include the need for obtaining large labeled datasets, high-quality features, and domain expertise in supervised learning. Deep learning, while not requiring feature extraction, still demands a substantial amount of labeled data and raises concerns about interpretability and overfitting (Cui et al., 2018).
Similarity-based link prediction methods continue to stay relevant, despite their usually lower prediction accuracy. This is because the size of graphs originated from the web, social networks or biological relations force us to use very simple algorithms if those graphs are to be computed in acceptable time (Garcia Gasulla and Cortés García, 2014). Similarity-based link prediction methods are highly cost-effective, as they offer competitive prediction quality with their low complexity in time and space (Zhou, 2021). Further, in certain applications like friend recommendation, preference is given to predictions with explanations, a feature not readily achievable through machine learning techniques (Barbieri et al., 2014).
Yang et al. (Yang et al., 2015b) introduce the Local Neighbors Link (LNL) measure, motivated by cohesion between common neighbors and predicted nodes, and implemented it on both MapReduce and Spark. Cui et al. (Cui et al., 2016) present a parallel algorithm for efficiently evaluating Common Neighbors (CN) similarity, obtaining node pairs with CN values surpassing a specified lower bound. Guo et al. (Guo et al., 2019) propose Common Neighbour Tightness (CNT), incorporating the aggregation degree of common neighbors by considering their proximity through local information and neighborhood tightness. Rafiee et al. (Rafiee et al., 2020) introduce Common Neighbors Degree Penalization (CNDP), which factors in clustering coefficient as a structural property and considers neighbors of shared neighbors. Mumin et al. (Mumin et al., 2022) contribute an algorithm combining common neighbors and node degree distribution to estimate link presence likelihood between two nodes based on local information.
@@
As mentioned earlier, similarity-based algorithms are competitive to other high-quality dimensionality reduction and machine learning techniques, thanks to their simplicity, interpretability (Pai et al., 2019; Barbieri et al., 2014), and computational efficiency (Garcia Gasulla and Cortés García, 2014) — and are thus often combined with other techniques (Kumari et al., 2022; AbuOda et al., 2020; Pai et al., 2019). However, the evaluation of these algorithms on large networks is crucial, as testing on small networks can yield misleading conclusions (Zhou, 2021; Zhou et al., 2021). Despite this, a significant portion of the discussed works focuses on small (Guo et al., 2019; Rafiee et al., 2020; Mumin et al., 2022; Papadimitriou et al., 2012; Vega-Oliveros et al., 2021; Saifi et al., 2023; Ferreira et al., 2019; Benhidour et al., 2022) to medium-scale graphs (Yang et al., 2015b; Cui et al., 2016; Kalkan and Hambiralovic, [n. d.]; Mohan et al., 2017; Wang et al., 2019; Bastami et al., 2019; Shin et al., 2012; Garcia Gasulla and Cortés García, 2014), with less than a million or billion edges. Parallelism becomes essential on large networks, and while some approaches, such as the ones based on common neighbors (Yang et al., 2015b; Cui et al., 2016), random walks (Papadimitriou et al., 2012), community structures (Mohan et al., 2017; Wang et al., 2019), and approximation (Garcia Gasulla and Cortés García, 2014), incorporate parallelism, the design of suitable data structures for efficient score computation remains an often-overlooked aspect. This technical report aims to bridge both gaps, while also proposing a heuristic for efficient computation.
@@ 3. Preliminaries
In an undirected graph $G(V,E)$ with sets of vertices $V$ and edges $E$, link prediction aims to identify missing or future edges from the set $U-E$, where $U$ contains all possible $|V|(|V|-1)/2$ edges in the graph (Zhou, 2021). Link prediction, thus, is akin to finding a needle in a haystack, as the correct edges need to be identified within a vast set of incorrect ones (Garcia Gasulla and Cortés García, 2014; Wang et al., 2014). Garcia et. al (Garcia Gasulla and Cortés García, 2014) observe that this ratio goes from $1:11k$ in their best case to $1:27M$ in the worst case. Further, larger graphs more likely to be incomplete. These challenges make the study of link prediction crucial.
Link prediction often relies on similarity metrics between node pairs, reflecting the likelihood of missing or future links (Wang et al., 2014; Arrar et al., 2023). The idea is grounded in the tendency for users to connect with similar individuals. More similarity thus suggests a higher probability of a future link. A ranked list of potential links based on similarity scores can then be used to predict the top-$k$ links are most likely to appear (on were missing) (Wang et al., 2014). Similarity measures are commonly categorized into local, quasi-local, and global measures. Local / neighborhood-based metrics, like Common Neighbours (CN), are calculated based on neighborhood information within a path distance of two. Global indices use network-wide information, while quasi-local indices combine both for distances up to two (Arrar et al., 2023).
@@ 3.3. Baseline Approach for Neighborhood-based Link Prediction
A baseline approach for neighborhood-based link prediction involves computing similarity scores between all non-connected node pairs $\{(a,c)\ |\ a,c\in V;(a,c)\notin E\}$, and returning the node pairs with top-$k$ scores as the predicted links. The similarity score is computed by assessing the commonality of neighbors between the two nodes. This approach is used by an number of studies previously mentioned (Gatadi and Swarupa Rani, 2023; Saifi et al., 2023; Benhidour et al., 2022; Mumin et al., 2022; Rafiee et al., 2020; Guo et al., 2019; Yang et al., 2015b; Papadimitriou et al., 2012; Wang et al., 2019). Notably, popular network analysis software packages such as NetworKit (Staudt et al., 2016) and igraph (Csardi et al., 2006) also implement this baseline approach.
@@ 4.1.1. Avoiding node pairs with no common neighbors
As noted earlier, there are a vast number of node pairs $(a,c)$ in a graph with no common neighbors, i.e., $\Gamma_{a}\cap\Gamma_{c}=\phi$ (here $\Gamma_{a}$ represents the neighbors of vertex $a$), and thus have a similarity score of 00. This is especially true for large sparse graphs. To address this, we limit computation of similarity scores of each vertex $a$ in the graph to its second order neighbors $c$, i.e., to vertices that are neighbors of the immediate neighbors $a$. We do this with a Depth First Search (DFS) traversal from each vertex $a$, limited to a distance of $2$.
@@ 4.1.3. Design of Hashtable
As C++’s inbuilt map has poor performance, we employ collision-free per-thread hashtables (we discuss later about parallelization), similar to our previous work (Sahu, 2023). An example of this hashtable for two threads is shown in Figure 2. Each hashtable includes a keys vector, a values vector (of size $|V|$), and a key count. The value associated with each key is stored or accumulated in the index pointed to by the key. To prevent false cache sharing, we independently update the key count of each hashtable and allocate it separately on the heap. As previously mentioned, the vertex IDs of second-order neighbors $c$ of each vertex $a$ serve as keys in the hashtable, and the associated values represent the number of times $c$ has been visited during a DFS traversal (of distance $2$) from $a$. For AA and RA scores, we use a hashtable with integral keys and floating-point values.
(a) Our Improved Baseline approach (IBase)
(b) Disregard hubs with degree $>8$ (DLH)
(c) Disregard hubs with degree $>4$ (DLH)
Figure 1. Illustration of our link prediction approach which Disregards Large Hubs (DLH), i.e., $1^{st}$ order neighbors with high degree. Here we focus on the neighborhood of vertex $1$, but the approach applies to each vertex in the graph. In the figure, $1$ is outlined in black, its $1^{st}$ order neighbors in red, its $2^{nd}$ order neighbors in blue, and explored/traversed vertices are shown with a yellow fill. Edge directions indicate traversal, with some $2^{nd}$ order vertices omitted for simplicity (dotted edges). (a) Depicts the our Improved Baseline (IBase) approach, which considers all $2^{nd}$ order neighbors of vertex $1$. (b) Presents our DLH approach, which considers only $2^{nd}$ order neighbors linked to $1$ through a small hub (degree $\leq 8$). This pruning reduces runtime and enhances prediction quality. (c) Illustrates DLH approach, where vertices with degree $>4$ are considered large hubs.
Figure 2. Illustration of collision-free per-thread hashtables that are well separated in their memory addresses, for two threads. Each hashtable consists of a keys vector, values vector (of size $|V|$), and a key count ($N_{0}$/$N_{1}$). The value associated with each key is stored/accumulated in the index pointed by the key. As the key count of each hashtable is updated independently, we allocate it separately on the heap to avoid false cache sharing (Sahu, 2023). These are used in the scoring phase of our implementation to track common neighbors.
4.1.4. Avoiding first order neighbors
@@ 4.1.6. Tracking top-k edges
Since storing all obtained similarity scores is impractical due to the scale of processed graphs, we use a min-heap based prediction list. This min-heap enables us to maintain the top-$k$ edges with the highest similarity scores per thread. It works by evicting the node pair with the lowest score when a new node pair with a higher score is encountered. As an optimization, we convert the per-thread prediction lists into a min-heap only after its have been populated with $k$ entries.
@@ 4.1.7. Parallelizing the process
4.2. Our Heuristic Approach which Disregards Large Hubs (DLH)
@@ 4.2.1. Explanation of the approach
To optimize link prediction further, we recognize that low-degree nodes, representing users with few connections in the social network, are more selective in accepting friend requests and are likely to form connections with people they have stronger, more meaningful relationships with, such as close friends and family. Thus, low-degree nodes confer significant similarity among their neighbors, while high-degree nodes generally do not (due to their lack of selectivity). We call such high-degree nodes as large hubs. Accordingly, for a given vertex $a$ and each of its immediate neighbors $b$, we only explore the neighbors $c\in\Gamma_{b}$ of $b$ if the degree of $b$ is within a certain threshold $L_{H}$, i.e., $|\Gamma_{b}|\leq L_{H}$. We call this threshold as the hub limit. This minimizes DFS exploration of second-degree neighbors $c$ during computation of neighbor-based similarity scores.

### v4

@@ 1. Introduction
Most real-world networks are incomplete (Kim and Leskovec, 2011; Wang et al., 2014). These networks lie somewhere in the range of a deterministic and a purely random structure, and are thus partially predictable (Lü et al., 2015). Link prediction is the problem of identifying potentially missing/undiscovered connections in such networks (Marchette and Priebe, 2008; Kim and Leskovec, 2011), or even forecasting future connections (Bringmann et al., 2010; Juszczyszyn et al., 2011), by examining the current network structure. This is useful in various applications, such as recommending items for online purchase (Akcora et al., 2011), helping people to find potential collaborators (Mori et al., 2012; Tang et al., 2012), assessing the trustworthiness of individuals (Alnumay et al., 2019), uncovering criminal activities and individuals (Berlusconi et al., 2016; Lim et al., 2019), and predicting new protein-protein interactions or generating hypotheses (Cannistraci et al., 2013; Nasiri et al., 2021).
@@
Similarity measures are frequently employed to predict the likelihood of missing or future links between unconnected nodes in a network (Wang et al., 2014; Arrar et al., 2023). The principle is straightforward: higher similarity indicates a greater likelihood of connection (Wang et al., 2014). The choice of metric depends on the network’s characteristics, with no single metric dominating across different datasets (Arrar et al., 2023; Zhou, 2021). Local / neighborhood-based similarity metrics such as Common Neighbors, Jaccard Coefficient, Sørensen Index, Salton Cosine similarity, Hub Promoted, Hub Depressed, Leicht-Holme-Nerman, Adamic-Adar, and Resource Allocation, which are based on neighborhood information within a path distance of two, remain popular (Arrar et al., 2023; Wang et al., 2014). This is due to their simplicity, interpretability (Pai et al., 2019; Barbieri et al., 2014), computational efficiency (Garcia Gasulla and Cortés García, 2014), and the ability to capture underlying structural patterns.
@@
However, many studies (Gatadi and Swarupa Rani, 2023; Saifi et al., 2023; Benhidour et al., 2022; Mumin et al., 2022; Rafiee et al., 2020; Guo et al., 2019; Yang et al., 2015b; Papadimitriou et al., 2012; Wang et al., 2019) and network analysis software (Staudt et al., 2016; Csardi et al., 2006) use a baseline approach for link prediction, computing unnecessary similarity scores for all non-connected node pairs. Further, early studies often evaluate a limited number of algorithms on small networks — this can result in misleading conclusions (Zhou, 2021; Zhou et al., 2021). Further, much-existing research does not address link prediction for large networks with close to a billion edges (Muscoloni et al., 2022; Mumin et al., 2022; Nasiri et al., 2021; Xian et al., 2021; Ghasemian et al., 2020; Mara et al., 2020; Wang et al., 2019; Xu et al., 2019; Mohan et al., 2017; Cui et al., 2016; Garcia Gasulla and Cortés García, 2014; Papadimitriou et al., 2012; Wang et al., 2023b, c, a). As the collection of data, represented as graphs, reaches unprecedented levels, it becomes necessary to design efficient parallel algorithms for link prediction on such graphs. While link prediction algorithms are often pleasingly parallel, most studies do not address the design of suitable data structures for efficient computation of scores.
@@
Further, the link prediction problem faces significant imbalance, with the number of known links often being several orders of magnitude less than non-existent links. This imbalance hinders the effectiveness of many link prediction methods, particularly on large networks (Wang et al., 2014; Garcia Gasulla and Cortés García, 2014). Thus, heuristics are needed to minimize the computation needed, without sacrificing on quality. Quality assessment measures for link prediction include precision, recall, F1 score, accuracy, and Area Under the Receiver Curve (AUC). While AUC is commonly used (Arrar et al., 2023), it may provide misleading results (Yang et al., 2015a; Lichtnwalter and Chawla, 2012), leading to our focus on the F1 score.
@@ 2. Related work
While effective in capturing nonlinear relationships for accurate predictions in complex networks, embedding-based methods face challenges with high-centrality nodes. Matrix factorization methods often struggle with complex network structures, require substantial computational resources and pose overfitting risks (Arrar et al., 2023; Martínez et al., 2016). Machine learning based approaches have their own set of drawbacks. These include the need to obtain large labeled datasets, high-quality features, and domain expertise in supervised learning. Deep learning, while not requiring feature extraction, still demands a substantial amount of labeled data and raises concerns about interpretability and overfitting (Cui et al., 2018).
Similarity-based link prediction methods continue to stay relevant, despite their usually lower prediction accuracy. This is because the size of graphs originating from the web, social networks or biological relations force us to use very simple algorithms if those graphs are to be computed in acceptable time (Garcia Gasulla and Cortés García, 2014). Similarity-based link prediction methods are highly cost-effective, as they offer competitive prediction quality with their low complexity in time and space (Zhou, 2021). Further, in certain applications like friend recommendation, preference is given to predictions with explanations, a feature not readily achievable through machine learning techniques (Barbieri et al., 2014).
Yang et al. (Yang et al., 2015b) introduce the Local Neighbors Link (LNL) measure, motivated by cohesion between common neighbors and predicted nodes, and implemented it on both MapReduce and Spark. Cui et al. (Cui et al., 2016) present a parallel algorithm for efficiently evaluating Common Neighbors (CN) similarity, obtaining node pairs with CN values surpassing a specified lower bound. Guo et al. (Guo et al., 2019) propose Common Neighbour Tightness (CNT), incorporating the aggregation degree of common neighbors by considering their proximity through local information and neighborhood tightness. Rafiee et al. (Rafiee et al., 2020) introduce Common Neighbors Degree Penalization (CNDP), which factors in the clustering coefficient as a structural property and considers neighbors of shared neighbors. Mumin et al. (Mumin et al., 2022) contribute an algorithm combining common neighbors and node degree distribution to estimate link presence likelihood between two nodes based on local information.
@@
As mentioned earlier, similarity-based algorithms are competitive with other high-quality dimensionality reduction and machine learning techniques, thanks to their simplicity, interpretability (Pai et al., 2019; Barbieri et al., 2014), and computational efficiency (Garcia Gasulla and Cortés García, 2014) — and are thus often combined with other techniques (Kumari et al., 2022; AbuOda et al., 2020; Pai et al., 2019). However, the evaluation of these algorithms on large networks is crucial, as testing on small networks can yield misleading conclusions (Zhou, 2021; Zhou et al., 2021). Despite this, a significant portion of the discussed works focuses on small (Guo et al., 2019; Rafiee et al., 2020; Mumin et al., 2022; Papadimitriou et al., 2012; Vega-Oliveros et al., 2021; Saifi et al., 2023; Ferreira et al., 2019; Benhidour et al., 2022; Wang et al., 2023b, c, a) to medium-scale graphs (Yang et al., 2015b; Cui et al., 2016; Kalkan and Hambiralovic, [n. d.]; Mohan et al., 2017; Wang et al., 2019; Bastami et al., 2019; Shin et al., 2012; Garcia Gasulla and Cortés García, 2014), with less than a million or billion edges. Parallelism becomes essential on large networks, and while some approaches, such as the ones based on common neighbors (Yang et al., 2015b; Cui et al., 2016), random walks (Papadimitriou et al., 2012), community structures (Mohan et al., 2017; Wang et al., 2019), and approximation (Garcia Gasulla and Cortés García, 2014), incorporate parallelism, the design of suitable data structures for efficient score computation remains an often-overlooked aspect. This technical report aims to bridge both gaps, while proposing a heuristic for efficient computation.
@@ 3. Preliminaries
In an undirected graph $G(V,E)$ with sets of vertices $V$ and edges $E$, link prediction aims to identify missing or future edges from the set $U-E$, where $U$ contains all possible $|V|(|V|-1)/2$ edges in the graph (Zhou, 2021). Link prediction, thus, is akin to finding a needle in a haystack, as the correct edges need to be identified within a vast set of incorrect ones (Garcia Gasulla and Cortés García, 2014; Wang et al., 2014). Garcia et. al (Garcia Gasulla and Cortés García, 2014) observe that this ratio goes from $1:11k$ in their best case to $1:27M$ in the worst case. Further, larger graphs are more likely to be incomplete. These challenges make the study of link prediction crucial.
Link prediction often relies on similarity metrics between node pairs, reflecting the likelihood of missing or future links (Wang et al., 2014; Arrar et al., 2023). The idea is grounded in the tendency for users to connect with similar individuals. More similarity thus suggests a higher probability of a future link. A ranked list of potential links, based on similarity scores, can then be used to predict the top-$k$ links that are most likely to appear (or were missing) (Wang et al., 2014). Similarity measures are commonly categorized into local, quasi-local, and global measures. Local / neighborhood-based metrics, like Common Neighbours (CN), are calculated based on neighborhood information within a path distance of two. Global indices use network-wide information, while quasi-local indices combine both for distances up to two (Arrar et al., 2023).
@@ 3.3. Baseline approach for Neighborhood-based Link Prediction
A baseline approach for neighborhood-based link prediction involves computing similarity scores between all non-connected node pairs $\{(a,c)\ |\ a,c\in V;(a,c)\notin E\}$, and returning the node pairs with top-$k$ scores as the predicted links. The similarity score is computed by assessing the commonality of neighbors between the two nodes. This approach is used by a number of studies previously mentioned (Gatadi and Swarupa Rani, 2023; Saifi et al., 2023; Benhidour et al., 2022; Mumin et al., 2022; Rafiee et al., 2020; Guo et al., 2019; Yang et al., 2015b; Papadimitriou et al., 2012; Wang et al., 2019). Notably, popular network analysis software packages such as NetworKit (Staudt et al., 2016) and igraph (Csardi et al., 2006) also implement this baseline approach.
@@ 4.1.1. Avoiding node pairs with no common neighbors
As noted earlier, there are a vast number of node pairs $(a,c)$ in a graph with no common neighbors, i.e., $\Gamma_{a}\cap\Gamma_{c}=\phi$ (here $\Gamma_{a}$ represents the neighbors of vertex $a$), and thus have a similarity score of 00. This is especially true for large, sparse graphs. To address this, we limit the computation of similarity scores of each vertex $a$ in the graph to its second-order neighbors $c$, i.e., to vertices that are neighbors of the immediate neighbors $a$. We do this with a Depth First Search (DFS) traversal from each vertex $a$, limited to a distance of $2$.
@@ 4.1.3. Design of Hashtable
As C++’s inbuilt map has poor performance, we employ collision-free per-thread hashtables (we discuss parallelization later), similar to our previous work (Sahu, 2023). An example of this hashtable for two threads is shown in Figure 1. Each hashtable includes a keys vector, a values vector (of size $|V|$), and a key count. The value associated with each key is stored or accumulated in the index pointed to by the key. To prevent false cache sharing, we independently update the key count of each hashtable and allocate it separately on the heap. As previously mentioned, the vertex IDs of second-order neighbors $c$ of each vertex $a$ serve as keys in the hashtable, and the associated values represent the number of times $c$ has been visited during a DFS traversal (of distance $2$) from $a$. For AA and RA scores, we use a hashtable with integral keys and floating-point values.
Figure 1. Illustration of collision-free per-thread hashtables that are well separated in their memory addresses, for two threads. Each hashtable consists of a keys vector, a values vector (of size $|V|$), and a key count ($N_{0}$/$N_{1}$). The value associated with each key is stored/accumulated in the index pointed by the key. As the key count of each hashtable is updated independently, we allocate it separately on the heap to avoid false cache sharing (Sahu, 2023). These are used in the scoring phase of our implementation to track common neighbors.
4.1.4. Avoiding first-order neighbors
@@ 4.1.6. Tracking top-k edges
Since storing all obtained similarity scores is impractical due to the scale of processed graphs, we use a min-heap based prediction list. This min-heap enables us to maintain the top-$k$ edges with the highest similarity scores per thread. It works by evicting the node pair with the lowest score when a new node pair with a higher score is encountered. As an optimization, we convert the per-thread prediction lists into a min-heap only after they have been populated with $k$ entries.
@@ 4.1.7. Parallelizing the process
4.2. Our Disregard Large Hubs (DLH) heuristic
@@ 4.2.1. Explanation of the approach
To optimize link prediction further, we recognize that low-degree nodes, representing users with few connections in the social network, are more selective in accepting friend requests and are likely to form connections with people they have stronger, more meaningful relationships with, such as close friends and family. Thus, low-degree nodes confer significant similarity among their neighbors, while high-degree nodes generally do not (due to their lack of selectivity). We refer to such high-degree nodes as large hubs. Accordingly, for a given vertex $a$ and each of its immediate neighbors $b$, we only explore the neighbors $c\in\Gamma_{b}$ of $b$ if the degree of $b$ is within a certain threshold $L_{H}$, i.e., $|\Gamma_{b}|\leq L_{H}$. We call this threshold the hub limit. This minimizes DFS exploration of second-degree neighbors $c$ during the computation of neighbor-based similarity scores.